OpenStudy (anonymous):

meade gave 1/4 of candy to max and then ate 1/2 of candy bar himself. what fraciton of the candy bar is left?

OpenStudy (anonymous):

1/4 1-1/2-1/4 = 1/4

OpenStudy (anonymous):

if i am not mistaken if you gave away 1/4 of candy to max and then ate 1/2 then i believ it should be 1/4 for the answer because 4 quarters equals on full bar of candy

OpenStudy (anonymous):

sorry it is gave 1/4 and ate 1/3

OpenStudy (anonymous):

1 -1/4 - 1/3 = 12/12 - 3/12-4/12 = 5/12